The Local L^{A}TEX Support Strategy

Martin Reddy

Version 1.0 — September 4, 1996

Mission Statement: given the current support staff shortages, it is hoped to be able to maintain a functional L^{A}TEX distribution and support mechanism which meets the reasonable needs of local users, whilst remaining compatible with the standard distribution as found at other sites.

Executive Summary

For those who never read documentation, the following is a summary of the support levels for the various facets of the local L^{A}TEX installation. Those who actually use L^{A}TEX on a regular basis would be advised to acquaint themselves with the rest of this document so that they fully understand the local support situation.

Category  Supported

Communication Channels

latex@dcs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Partially
cs.latex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Partially

TEX/LaTeX installation upgrades

Unix platform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  No
Mac platform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  No

Packages & Classes

Local packages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  Yes
Standard L^{A}TEXpackages . . . . . . . . . . . . . . . . . . . . . . . . Partially
Non-standard packages . . . . . . . . . . . . . . . . . . . . . . . . . .  No

Documentation

Web pages (inc. local FAQ) . . . . . . . . . . . . . . . . . . . . .  Yes
Standard L^{A}TEX documentation . . . . . . . . . . . . . . . . . . Partially
Non-standard documentation . . . . . . . . . . . . . . . . . . . .  No

Fonts

Current font installation . . . . . . . . . . . . . . . . . . . . . . . . Partially
Additional fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  No

Auxiliary Utilities

Standard (CTAN) utilities . . . . . . . . . . . . . . . . . . . . . . Partially
Non-standard utilities . . . . . . . . . . . . . . . . . . . . . . . . . . .  No

To summarise further: local and standard packages will receive highest priority. Documentation will be kept reasonably current. Fonts and auxiliary utilities (e.g. BibTEX and dvi utilities) will receive lower priority. Non-standard material and non-Unix platforms will not be supported, although some assistance may be available.

1 The Current Situation

The current full CTAN distribution consists of some 60,000 files spread over 4,500 directories consuming well over 2 GB of disk space. This mass is provided ‘as-is’ with no standard mechanism to easily integrate new upgrades with the existing system.

The current departmental support for L^{A}TEX consists of one person, working part-time, with only a few hours a week to dedicate to the maintenance of the local system. It is therefore infeasible to provide full and immediate support for all aspects of our L^{A}TEX installation. As a result, this document has been prepared to give users a clear picture of the feasible support that can be offered for the various facets of our L^{A}TEX system.

2 The Support Strategy

2.1 Communication Channels

Mail sent to the latex@dcs account will receive an automatic reply stating that the department is unable to provide support for the L^{A}TEX package due to staff shortages. All mail will be stored in the latex account until the staff situation is resolved. However, the L^{A}TEX support person will occasionally check this mailbox to see if there are any relevant queries requiring action.

The L^{A}TEX support person will also periodically check the local cs.latex newsgroup and answer any questions where required.

2.2 L^{A}TEX Installation Updates

New versions of the L^{A}TEX 2ε and TEX systems are distributed fairly frequently. It will not be possible to ensure that the local system is always updated with the most recent versions. However, new versions will be installed should our local system prove to be substantially out of date, and if there is available time.

Support for L^{A}TEX updates on other platforms (e.g. OzTEX on the Macs) will not be available.

The old version of L^{A}TEX (version 2.09) and the old version of TEX (version C 3.141) will remain on the system, but will not be supported in any way.

2.3 Packages and Classes

Local L^{A}TEX 2ε packages and classes will receive high priority (e.g. csletter, csthesis, cstechrep etc.) Standard packages and classes available from CTAN will be kept reasonable up to date, although normally on a user demand basis.

Non standard L^{A}TEX 2ε packages and classes (i.e. those not found at CTAN), and old L^{A}TEX (2.09) style files will not be supported in any way.

2.4 Documentation

The primary source for locating documentation will be the local L^{A}TEX World Wide Web page at: http://www.dcs.ed.ac.uk/packages/latex. This will provide links to all on-line documentation and will be kept reasonably current. (This includes the local L^{A}TEX FAQ.)

Documentation for local packages and classes will be kept as up to date as possible. Documentation for any of the standard packages or classes will be kept reasonably current (where available). In all cases, documentation should always be in-line with their associated package or class versions.

Documentation for non standard material will not be supported.

2.5 Fonts

We currently have a vast selection of fonts already installed on the system. These include all of the standard fonts which can be found at most other sites. Any problems with the current suite of fonts will be addressed in due course, however these requests will receive a low priority.

New fonts will not be added to the system: users wishing to use any font which does not exist in the current installation must support this themselves in their own directory space, i.e. in their ∼/tex/fonts/... directory (with subdirectories tfm, vf, pk, type1, source).

2.6 Auxiliary Utilities

Standard utilities which are used by L^{A}TEX users or the L^{A}TEX system (such as the BibTEX tools, DVI utilities, PS utilities, kpathsea tools etc.) will be supported, but at a low priority.

Any non-standard utilities which individual users support, or make use of on a personal basis, will not be supported.